# A Simple Framework for Open-Vocabulary Zero-Shot Segmentation

### v2

@@ Abstract
Zero-shot classification capabilities naturally arise in models trained within a vision-language contrastive framework. Despite their classification prowess, these models struggle in dense tasks like zero-shot open-vocabulary segmentation. This deficiency is often attributed to the absence of localization cues in captions and the intertwined nature of the learning process, which encompasses both image representation learning and cross-modality alignment. To tackle these issues, we propose SimZSS, a Simple framework for open-vocabulary Zero-Shot Segmentation. The method is founded on two key principles: i) leveraging frozen vision-only models that exhibit spatial awareness while exclusively aligning the text encoder and ii) exploiting the discrete nature of text and linguistic knowledge to pinpoint local concepts within captions. By capitalizing on the quality of the visual representations, our method requires only image-caption pairs datasets and adapts to both small curated and large-scale noisy datasets. When trained on COCO Captions across 8 GPUs, SimZSS achieves state-of-the-art results on 7 out of 8 benchmark datasets in less than 15 minutes.
Figure 1: Visualization of the patch-level visual representations and text concepts in the RGB color space. PCA is used to map the dense representation of a single image into a three-dimensional space. The three-dimensional representations (color) of the text concepts from the concept bank and the corresponding caption are obtained and shown for each image. Each row includes the original image and dense feature visualization at different resolutions. These include the training resolution ($16\times 16$) and higher resolutions (2×2$\times$2 ×, 4×4$\times$4 ×, and 8×8$\times$8 ×).
@@ 1 Introduction
Semantic segmentation stands as a cornerstone task within the realm of computer vision, playing a pivotal role in various applications ranging from autonomous driving to medical image analysis. However, its widespread adoption and scalability are hindered by the inherent label-intensive nature of traditional methods, demanding copious amounts of fine-grained annotations for training. Furthermore, traditional approaches, e.g.,  Strudel et al. (2021); Cheng et al. (2022) are typically closed-vocabulary, meaning they only work for a pre-defined set of categories and generalize poorly to unseen classes. Self-supervised learning paradigms (Caron et al., 2021; Zhou et al., 2022b; He et al., 2020; Balazevic et al., 2024; Lebailly et al., 2023), which use pretext tasks to learn discriminative representations from data itself, offer a promising solution to alleviate the annotation burden. Representations obtained in this manner are typically clustered semantically, potentially even at a fine-grained level (Balažević et al., 2023; Lebailly et al., 2023; Stegmüller et al., 2023) and, as such, yield excellent performance in various applications. Nonetheless, a degree of labeling is still necessary, whether for finetuning or constructing the support set of a $k$-NN classifier, for instance.
@@
Recently, vision-language contrastive learning has proven to be a simple and effective approach for transforming web-scale amounts of noisy image-caption pairs into zero-shot classification capabilities (Radford et al., 2021; Jia et al., 2021). Nonetheless, the substantial computational and data requirements of these methods pose significant challenges. To address these limitations, LiT (Zhai et al., 2022) proposed aligning a text encoder with a pretrained and frozen pure vision model. By decoupling the image representation learning process from the vision-language alignment process, this methodology enhances both computational and data efficiency while also improving performance in zero-shot classification tasks. Orthogonal to this, various studies (Hamilton et al., 2022; Siméoni et al., 2021; Wang et al., 2023; Zhang et al., 2023; Siméoni et al., 2023) have observed and capitalized on the spatial awareness of self-supervised pretrained vision transformers (ViTs). These works have demonstrated remarkable capabilities for dense downstream tasks with minimal or even zero trainable parameters (Balazevic et al., 2024). This evidence suggests that self-supervised vision transformers may just be one text encoder away from becoming open-vocabulary segmenters. We investigate this hypothesis and propose a simple framework for open-vocabulary zero-shot segmentation. In essence, we leverage linguistic knowledge to identify concepts in captions and build upon the quality of a frozen vision tower to retrieve corresponding concepts in images.
@@ 2 Related works
Traditionally, computer vision methods operate under the closed-vocabulary hypothesis. This assumption presumes that all object categories a model is expected to classify, detect, or segment at test time are already known and labeled during training. This presents significant challenges due to the extensive labeling required and the limited generalizability of the resulting models. Open-vocabulary learning aims to eliminate these limitations. Notably, this is a more challenging objective than the one targeted by self-supervised visual representation learning (Chen & He, 2020; Caron et al., 2021; Li et al., 2021; Grill et al., 2020; Oquab et al., 2023), which assumes the availability of labels at test time. To address this additional constraint, a text encoder is typically trained jointly with the vision tower to maximize vision-language alignment on large amounts of image-caption pairs via contrastive learning (Radford et al., 2021; Jia et al., 2021). Subsequent studies have either focused on improving the cross-modality alignment pretext task (Alayrac et al., 2022; Yu et al., 2022; Tschannen et al., 2024) or on improving the computational efficiency (Zhai et al., 2022; Li et al., 2023). However, while these models excel at classification, their performance on dense downstream tasks is usually subpar.
@@
To overcome this limitation, approaches that do not require segmentation masks have been proposed. PACL (Mukhoti et al., 2023) demonstrates the effectiveness of training only the projections at the interface of the two modalities and bootstrapping patch-level cross-modal similarities. Similarly, TCL (Cha et al., 2023) leverages existing fine-grained similarities to jointly learn text-grounded masks and impose contrastive region-text alignment based on the obtained masks. Alternatively, GroupViT Xu et al. (2022) proposes a specialized architecture that performs hierarchical grouping/pooling of visual tokens based on their similarity with learnable query tokens. Another line of work showed the benefits of integrating a consistency objective between augmented and/or masked views of the input image (Dong et al., 2023; Ren et al., 2023) in the context of vision-language contrastive learning. Related to that, ReCo (Shin et al., 2022) reported improved localization capabilities by exploiting the co-occurrence of objects across multiple images. Finally, Wysoczańska et al. (2024; 2024); Rewatbowornwong et al. (2023) combine the vision-language alignment capabilities of CLIP (Radford et al., 2021) with the spatial-awareness of self-supervised vision transformers, e.g., DINO (Caron et al., 2021) to develop open-vocabulary zero-shot segmenter.
@@ 3.1 Preliminaries
Transformers encode input signals as a sequence of tokens, and we refer to the corresponding output sequence as the dense representation. For vision transformers, this corresponds to a tensor $\bm{z}_{v}\in\mathbb{R}^{n_{v}\times d_{v}}$, where $d_{v}$ is the dimension of the representation space and $n_{v}$ is the number of patches in the input image. For text input, the dense representation $\bm{z}_{t}\in\mathbb{R}^{n_{t}\times d_{t}}$ is a sequence of $d_{t}$-dimensional tokens. The number of tokens, $n_{t}$, is determined by the byte-pair encoding algorithm (Gage, 1994).
@@ 3.2.1 Textual concept identification
Given an image-caption pair $(\bm{x}_{v},\bm{x}_{t})$, we utilize the text modality to identify semantic concepts. Language adheres to structured grammatical rules, allowing us to leverage linguistic knowledge to pinpoint the main subjects and objects within a sentence. Since captions are intended to describe images, the primary subjects and objects in the captions are likely to correspond to visual elements in the images. To achieve this, we employ a part-of-speech (POS) tagger, which automatically classifies words according to their grammatical roles. In our context, noun phrases (NPs) are of particular interest, as they generally encapsulate the main subjects and objects in the sentences, providing a direct link to the visual concepts depicted in the images. However, captions can be quite noisy, so further filtering is necessary to discard NPs that are unlikely to appear in the images. To address this, we refine a noun phrase into a concept by restricting it to the noun and its first compound. We then discard any concept that is absent in a predefined bank of concepts.
@@ 3.2.2 Visual concept identification
where $\tau$ is a temperature parameter that regulates the sharpness of the similarity (see Tab. 7). Consequently, the $l^{th}$ visual concept is obtained by bootstrapping existing cross-modality similarities, i.e., via similarity-based pooling (see Fig. 2):
@@ 3.3 Cross-modality consistency
In Section 3.2, we outline a methodology for identifying pairs of local concepts across textual and visual modalities. This method leverages established similarities and thus may benefit from additional supervision at the beginning of training, as discussed in Section 3.3.1.
@@ 3.3.2 Concept-level consistency
After obtaining pairs of vision-language concepts, an intuitive approach to enforce consistency at the concept level is to use a contrastive objective, akin to the one described in Section 3.3.1, but applied between concepts. Empirically, we found that this approach did not yield the desired performance improvements on dense downstream tasks. At a global level, images and captions represent complex scenes, supporting the hypothesis that only $b$ positive image-caption pairs exist among the $b^{2}$ possible ones in a batch. Conversely, concepts typically encode individual objects that are likely to occur multiple times within a batch. This suggests that an instance-level objective is ill-suited for our setting. Therefore, we opt for a semantic-level objective.
@@
After $l2$-normalizing the columns of both $\mathbf{h}$ and $\mathbf{C}_{v}$, we project one onto the other to derive a probability distribution over the visual concepts:
[…]
@@
where $\lambda$ is a weighting parameter whose impact is ablated in Table 7.
@@ 4 Experiments
In this section, we investigate the properties of SimZSS through various experiments. Details of the experimental setup are provided in Appendix A, and additional experiments can be found in Appendix B.
@@ 4.1 Zero-shot segmentation of foreground
LiT (Zhai et al., 2022) (ViT-B + LAION-400M) 94M 63M 80.5 31.8 23.3 24.7 18.7
LiT (Zhai et al., 2022) (ViT-B + COCO Captions) 94M 63M 86.1 35.5 25.6 25.8 18.1
CLIP-DNOiser (Wysoczańska et al., 2024) - - 80.9 35.9 24.6 31.7 20.0
Ours
SimZSS (ViT-B + LAION-400M) 94M 63M 85.1 34.2 24.9 27.8 19.6
SimZSS (ViT-S + COCO Captions) 21M 40M 87.2 37.3 23.8 29.2 17.9
SimZSS (ViT-B + COCO Captions) 94M 63M 90.3 43.1 29.0 33.0 21.8
@@
We validate the proposed method on a pixel-level zero-shot segmentation task. In this scenario, the model relies solely on textual class descriptions to classify image pixels. However, accurately representing the background class can be challenging due to dataset-specific properties not fully captured by text descriptions. As such, we follow previous works (Wysoczańska et al., 2024; Cha et al., 2023) and first evaluate without the background class.
@@
Figure 3: Vision-language alignment of text concepts and dense visual representations. Concepts present in the image are embedded independently by the text encoder and then projected onto the representations of each patch within the image. The images are processed at a resolution of $896\times 896$ pixels, corresponding to 4×4$\times$4 × the training resolution. The alignment is performed on LAION-400M using a ViT-B/14 as the vision tower.
@@
We follow the MMSegmentation (Contributors, 2020) implementation of Cha et al. (2023). Specifically, images are resized to have a shorter side of 448 pixels, and inference is performed with a sliding window using a stride of 224 pixels. ImageNet templates from Radford et al. (2021) are used to contextualize each class name before obtaining their textual representation. Finally, the predictions obtained by projecting patch representations onto class names, are up-sampled using bilinear interpolation to restore the predictions to the original image size. We report the mIoU scores across five standard datasets, namely Pascal VOC (Everingham et al., 2012), Pascal Context (Mottaghi et al., 2014), COCO-Stuff (Caesar et al., 2018), Cityscapes (Cordts et al., 2016) and ADE20K (Zhou et al., 2017). For datasets that encompass the background class, the corresponding annotations and pixels are simply ignored. For the sake of clarity, we report the results without any post-processing techniques such as Pixel-Adaptive Mask Refinement (PAMR) (Araslanov & Roth, 2020).
@@
In Table 1, we observe that our method using a ViT-S/14 as the vision tower is already competitive and that shifting to a larger vision backbone yields state-of-the-art results on all datasets when trained on COCO Captions. In particular, it can be seen that the proposed pipeline significantly outperforms models trained under the same conditions with LiT (Zhai et al., 2022). This suggests that the reported performance is not solely attributable to the freezing of the vision tower or DINOv2 (Oquab et al., 2023) visual pretraining. The results obtained with LAION-400M indicate that SimZSS is compatible with large-scale uncurated datasets, exhibiting overall strong performance and improvements compared to models trained with LiT in the same setting. However, it is also apparent that both LiT and SimZSS seem to benefit more from curation than scale in the segmentation tasks. This trend contrasts with findings in classification tasks, as shown in Table 3.

### v3

@@ Abstract
Zero-shot classification capabilities naturally arise in models trained within a vision-language contrastive framework. Despite their classification prowess, these models struggle in dense tasks like zero-shot open-vocabulary segmentation. This deficiency is often attributed to the absence of localization cues in captions and the intertwined nature of the learning process, which encompasses both image/text representation learning and cross-modality alignment. To tackle these issues, we propose SimZSS, a Simple framework for open-vocabulary Zero-Shot Segmentation. The method is founded on two key principles: i) leveraging frozen vision-only models that exhibit spatial awareness while exclusively aligning the text encoder and ii) exploiting the discrete nature of text and linguistic knowledge to pinpoint local concepts within captions. By capitalizing on the quality of the visual representations, our method requires only image-caption pair datasets and adapts to both small curated and large-scale noisy datasets. When trained on COCO Captions across 8 GPUs, SimZSS achieves state-of-the-art results on 7 out of 8 benchmark datasets in less than 15 minutes. Our code and pretrained models are publicly available at https://github.com/tileb1/simzss.
Figure 1: Visualization of the patch-level representations and text concepts in the RGB color space. PCA is used to map the dense representation of a single image into a three-dimensional space. The three-dimensional representations (color) of the text concepts from the concept bank and the corresponding caption are obtained and shown for each image. Each row includes the original image and dense feature visualization at different resolutions. These include the training resolution ($16\times 16$) and higher resolutions ($2\times$, $4\times$, and $8\times$).
@@ 1 Introduction
Semantic segmentation stands as a cornerstone task within the realm of computer vision, playing a pivotal role in various applications ranging from autonomous driving to medical image analysis. However, its widespread adoption and scalability are hindered by the inherent label-intensive nature of traditional methods, demanding copious amounts of fine-grained annotations for training. Furthermore, traditional approaches, e.g.,  Strudel et al. (2021); Cheng et al. (2022) are typically closed-vocabulary, meaning they only work for a pre-defined set of categories and generalize poorly to unseen classes. Self-supervised learning paradigms (Caron et al., 2021; Zhou et al., 2022b; He et al., 2020; Balazevic et al., 2024; Lebailly et al., 2024), which use pretext tasks to learn discriminative representations from data itself, offer a promising solution to alleviate the annotation burden. Representations obtained in this manner are typically clustered semantically, potentially even at a fine-grained level (Balazevic et al., 2024; Lebailly et al., 2024; Stegmüller et al., 2023) and, as such, yield excellent performance in various applications. Nonetheless, a degree of labeling is still necessary, whether for finetuning or constructing the support set of a $k$-NN classifier, for instance.
@@
Recently, vision-language contrastive learning has proven to be a simple and effective approach for transforming web-scale amounts of noisy image-caption pairs into zero-shot classification capabilities (Radford et al., 2021; Jia et al., 2021). Nonetheless, the substantial computational and data requirements of these methods pose significant challenges. Furthermore, the performance of these methods on finer-grained downstream tasks, such as zero-shot segmentation tends to be subpar (see Tables˜1 and 2). We argue that this is not only the consequence of a lack of local vision-language alignment but also a hint that the learned visual representations therein might be suboptimal. Based on this same observation, LiT (Zhai et al., 2022) proposed aligning a text encoder with a pretrained and frozen pure vision model. By decoupling the image representation learning process from the vision-language alignment process, this methodology enhances both computational and data efficiency while also improving performance in zero-shot classification tasks. The strong performance of LiT suggests that achieving zero-shot classification or segmentation relies on two key requirements:
  1. 1.
    The vision encoder must be capable of clustering images or groups of pixels with shared semantics.
  2. 2.
    The text encoder must be sufficiently aligned with the visual modality to accurately label these clusters.
The moderate performance of frozen vision encoder pretrained with CLIP on pure vision tasks (see Section˜A.3) combined with the stellar results of LiT support that the second requirement is significantly easier to accomplish than the first one and that the first requirement can be addressed independently of the second. Fortunately, various studies (Hamilton et al., 2022; Siméoni et al., 2021; Wang et al., 2023; Zhang et al., 2023; Siméoni et al., 2023) have observed and capitalized on the spatial awareness of self-supervised pretrained vision transformers (ViTs). These works have demonstrated remarkable capabilities for dense downstream tasks with minimal or even zero trainable parameters (Balazevic et al., 2024). This evidence suggests that self-supervised vision transformers may just be one text encoder away from becoming open-vocabulary segmenters. We investigate this hypothesis and propose a simple framework for open-vocabulary zero-shot segmentation. In essence, we capitalize on a pretrained and frozen vision encoder that exhibits excellent fine-grained semantic coherence and focus solely on aligning a text encoder to it. To effectively enforce a localized consistency objective, we leverage linguistic knowledge to identify concepts in captions and build on the quality of the frozen vision encoder to retrieve the corresponding concepts in images.
@@ 2 Related works
Traditionally, computer vision methods operate under the closed-vocabulary hypothesis. This assumption presumes that all object categories a model is expected to classify, detect, or segment at the test time are already known and labeled during training. This presents significant challenges due to the extensive labeling required and the limited generalizability of the resulting models. Open-vocabulary learning aims to eliminate these limitations. Notably, this is a more challenging objective than the one targeted by self-supervised visual representation learning (Chen & He, 2021; Caron et al., 2021; Li et al., 2021; Grill et al., 2020; Oquab et al., 2023), which assumes the availability of labels at test time. To address this additional constraint, a text encoder is typically trained jointly with the vision tower to maximize vision-language alignment on large amounts of image-caption pairs via contrastive learning (Radford et al., 2021; Jia et al., 2021). Subsequent studies have either focused on improving the cross-modality alignment pretext task (Alayrac et al., 2022; Yu et al., 2022; Tschannen et al., 2024) or on improving the computational efficiency (Zhai et al., 2022; Li et al., 2023b). However, while these models excel at classification, their performance on dense downstream tasks is usually subpar.
@@
To overcome this limitation, approaches that do not require segmentation masks have been proposed. PACL (Mukhoti et al., 2023) demonstrates the effectiveness of training only the projections at the interface of the two modalities and bootstrapping patch-level cross-modal similarities. Similarly, TCL (Cha et al., 2023) leverages existing fine-grained similarities to jointly learn text-grounded masks and impose contrastive region-text alignment based on the obtained masks. Alternatively, GroupViT Xu et al. (2022) proposes a specialized architecture that performs hierarchical grouping/pooling of visual tokens based on their similarity with learnable query tokens. Another line of work showed the benefits of integrating a consistency objective between augmented and/or masked views of the input image (Dong et al., 2023; Ren et al., 2023) in the context of vision-language contrastive learning. Recent works (Shin et al., 2022; Sun et al., 2024) also showed the possibility of achieving excellent zero-shot segmentation capabilities without any additional training. Finally, Wysoczańska et al. (2024; 2023); Rewatbowornwong et al. (2023) combine the vision-language alignment capabilities of CLIP (Radford et al., 2021) with the spatial-awareness of self-supervised vision transformers, e.g., DINO (Caron et al., 2021) to develop open-vocabulary zero-shot segmenter.
@@ 3.1 Preliminaries
Transformers encode an input signal as a sequence of tokens, and we refer to the corresponding output sequence as the dense representation. For vision transformers, this corresponds to a tensor $\bm{z}_{v}\in\mathbb{R}^{n_{v}\times d_{v}}$, where $d_{v}$ is the dimension of the representation space and $n_{v}$ is the number of patches in the input image. For text input, the dense representation $\bm{z}_{t}\in\mathbb{R}^{n_{t}\times d_{t}}$ is a sequence of $d_{t}$-dimensional tokens.
@@ 3.2.1 Textual concept identification
Given an image-caption pair $(\bm{x}_{v},\bm{x}_{t})$, we utilize the text modality to identify semantic concepts. Language adheres to structured grammatical rules, allowing us to leverage linguistic knowledge to pinpoint the main subjects and objects within a sentence. Since captions are intended to describe images, the primary subjects and objects in the captions are likely to correspond to visual elements in the images. Therefore, we employ a part-of-speech (POS) tagger, which automatically classifies words according to their grammatical roles. In our context, noun phrases (NPs) are of particular interest, as they generally encapsulate the main subjects and objects in the sentences, providing a direct link to the visual concepts depicted in the images. However, captions can be quite noisy, so further filtering is necessary to discard NPs that are unlikely to appear in the images. To address this, we refine a noun phrase into a concept by restricting it to the noun and its first compound. We then discard any concept that is absent in a predefined bank of concepts.
@@ 3.2.2 Visual concept identification
where $\tau$ is a temperature parameter that regulates the sharpness of the similarity distribution (see Tab.˜9). Consequently, the $l^{th}$ visual concept is obtained by bootstrapping existing cross-modality similarities, i.e., via similarity-based pooling (see Fig.˜2):
@@ 3.3 Cross-modality consistency
In Section˜3.2, we outline a methodology for identifying pairs of local concepts across textual and visual modalities. This method leverages existing similarities and thus may benefit from additional supervision at the beginning of training, as discussed in Section˜3.3.1.
@@ 3.3.2 Concept-level consistency
After obtaining pairs of vision-language concepts, an intuitive approach to enforce consistency at the concept level is to use a contrastive objective, akin to the one described in Section˜3.3.1, but applied between concepts from each modality. Empirically, we found that this approach did not yield the desired performance improvements on dense downstream tasks. At a global level, images and captions represent complex scenes, supporting the hypothesis that only $b$ positive image-caption pairs exist among the $b^{2}$ possible ones in a batch. Conversely, concepts typically encode individual objects that are likely to occur multiple times within a batch. This suggests that an instance-level objective is ill-suited for our setting. Therefore, we opt for a semantic-level objective.
@@
After $l2$-normalizing the columns of both $\mathbf{h}$ and $\mathbf{C}_{v}$, we project one onto the other to derive a probability distribution:
$\mathbf{p}=\underset{k}{\texttt{softmax}}\left(\mathbf{C}_{v}\mathbf{h}^{\top}\right)$ (8)
@@
where $\lambda$ is a weighting parameter whose impact is ablated in Table˜9.
@@ 4 Experiments
In this section, we investigate the properties of SimZSS through various experiments. Additional experiments can be found in Appendix˜A.
4.1 Experimental setup
4.1.1 Pretraining datasets
4.1.2 Networks architectures
4.1.3 Optimization
For COCO Captions, we conduct training over 4M processed samples ($\sim 6.6$ epochs) using a global batchsize of 16,384. We incorporate a warm-up strategy spanning 10% of the training steps, linearly ramping up the learning rate until it reaches its peak value, chosen from the set {8e-5, 3e-5, 8e-6, 3e-6}111scaled by batchsize / 256. Subsequently, we employ a cosine decay schedule for the remaining steps. Similarly, for LAION-400M, we train for 1 epoch with a global batchsize of 32,768, and we set the learning rate from the options {3e-5, 8e-6, 3e-6, 8e-7, 3e-7}. The remaining optimization settings align with those of OpenCLIP (Ilharco et al., 2021).
4.1.4 Text concepts
@@ 4.2 Zero-shot segmentation of foreground
CLIP-DINOiser (Wysoczańska et al., 2023) - - 80.9 35.9 24.6 31.7 20.0
LAION-400M
CLIP (Radford et al., 2021) (ViT-B) 0 157M 35.1 7.7 4.2 1.8 2.0
LiT (Zhai et al., 2022) (ViT-B) 94M 63M 80.5 31.8 23.3 24.7 18.7
SimZSS (ViT-B) 94M 63M 85.1 34.2 24.9 27.8 19.6
COCO Captions
LiT (Zhai et al., 2022) (ViT-B) 94M 63M 86.1 35.5 25.6 25.8 18.1
SimZSS (ViT-S) 21M 40M 87.2 37.3 23.8 29.2 17.9
SimZSS (ViT-B) 94M 63M 90.3 43.1 29.0 33.0 21.8
@@
We validate the proposed method on a pixel-level zero-shot segmentation task. In this scenario, the model relies solely on textual class descriptions to classify image pixels. However, accurately representing the background class can be challenging due to dataset-specific properties not fully captured by text descriptions. As such, we follow previous works (Wysoczańska et al., 2023; Cha et al., 2023) and first evaluate without the background class.
@@
Figure 3: Vision-language alignment of text concepts and dense visual representations. Concepts present in the image are embedded independently by the text encoder and then projected onto the representations of each patch within the image. The images are processed at a resolution of $896\times 896$ pixels, corresponding to $4\times$ the training resolution. The alignment is performed on LAION-400M using a ViT-B/14 as the vision tower.
@@
We follow the MMSegmentation (Contributors, 2020) implementation of Cha et al. (2023). Specifically, images are resized to have a shorter side of 448 pixels, and inference is performed with a sliding window of $448\times 448$ and a stride of 224 pixels. ImageNet templates from Radford et al. (2021) are used to contextualize each class name before obtaining their textual representation. Finally, the predictions obtained by projecting patch representations onto class names, are up-sampled using bilinear interpolation to restore the predictions to the original image size. We report the mIoU scores across five standard datasets, namely Pascal VOC (Everingham et al., 2012), Pascal Context (Mottaghi et al., 2014), COCO-Stuff (Caesar et al., 2018), Cityscapes (Cordts et al., 2016) and ADE20K (Zhou et al., 2017). For datasets that encompass the background class, the corresponding annotations and pixels are simply ignored. For the sake of clarity, we report the results without any post-processing techniques such as Pixel-Adaptive Mask Refinement (PAMR) (Araslanov & Roth, 2020).
@@
In Table˜1, we observe that our method using a ViT-S/14 as the vision tower is already competitive and that shifting to a larger vision backbone, namely a ViT-B/14, yields state-of-the-art results on all datasets when trained on COCO Captions. In particular, it can be seen that the proposed pipeline significantly outperforms models pretrained with LiT (Zhai et al., 2022) under the same conditions. This suggests that the reported performance is not solely attributable to the freezing of the vision tower or DINOv2 (Oquab et al., 2023) visual pretraining. The results obtained with LAION-400M indicate that SimZSS is compatible with large-scale uncurated datasets, exhibiting overall strong performance and improvements compared to models trained with LiT in the same setting. However, it is also apparent that both LiT and SimZSS seem to benefit more from curation than scale for segmentation. This trend contrasts with findings in classification tasks, as shown in Table˜3.
Finally, we present the performance of a CLIP-based model trained with a ViT-B/16 vision backbone on the LAION-400M dataset for 32 epochs (Ilharco et al., 2021). The results highlight that, while CLIP excels at image-level tasks such as zero-shot classification, it demonstrates limitations in achieving fine-grained cross-modal alignment. In Section˜A.3, we address this limitation and provide evidence suggesting that it is not solely attributable to vision-language misalignment.